# Weak continuous measurements require more work than strong ones

### v2

@@ III Thermodynamic analysis
We assume the ancilla is at time $t_{0}$ in a free state at equilibrium with its environment at inverse temperature $\beta$.
@@
$\text{W}_{\text{dr}}=\Delta E_{SA}=\Delta E_{S}+\Delta E_{A},$ (5)
where $\Delta E_{i}=E_{i}\left(t_{1}\right)-E_{i}\left(t_{0}\right)$, for $i=S,A$ represent the system’s (resp. ancilla’s) energy change. Above, $t_{1}$ is the final time of step $1$, while $t_{\text{f}}$ is the time at which the ancilla is readout, after the pure dephasing took place. This energy exchange arises from the dynamical evolution of the full-system state governed by the system-ancilla Hamiltonian. In Eq. (5), we have assumed that the coupling between the ancilla and the system vanishes at $t_{0}$ and at $t_{1}$. Moreover, we use that the pure dephasing does not alter the energy of the system and of the ancilla, such that the energy variations involved in Eq. (5) can also be computed between times $t_{0}$ and $t_{f}$, i.e. $\Delta E_{i}=E_{i}\left(t_{\text{f}}\right)-E_{i}\left(t_{0}\right)$.
@@
During the dephasing process, there is no energy exchange, neither work nor heat, with the systems; however, the total entropy increases by an amount:
$\sigma=\Delta S_{SAM}\geq 0.$ (6)
Additionally, the classical readout of the ancilla can, in principle, be implemented at no work cost [32].
@@
$W^{A}_{\text{reset}}\geq-\Delta E_{A}+\frac{1}{\beta}\Delta{\left\langle{S_{A}}\right\rangle},$ (7)
@@
represents the average change in entropy of the ancilla, realization-wise, expressed in terms of $\hat{\rho}_{A|r}=\text{Tr}_{S}\{\hat{\rho}_{SA|r}\}$. Based on the block-diagonal structure of $\hat{\rho}_{SA|r}$ in Eq. 4, the average entropy of the conditional ancilla states fulfills:
$\sum_{r}p_{r}S[\hat{\rho}_{A|r}(t_{f})]=H\left(p_{n}\right)+\sum_{n}p_{n}H\left(p_{r|n}\right)-H\left(p_{r}\right).$ (9)
@@
On the other hand, Landauer’s principle states that the work required to reset the classical memory fulfills [32]:
$W_{\text{reset}}^{M}\geq\frac{1}{\beta}H\left(p_{r}\right),$ (10)
where $H\left(p_{r}\right)=-\sum_{r}p_{r}\text{log}\left(p_{r}\right)$ is the Shanon entropy of probability distribution $p_{r}$ of obtaining measurement result $r$.
@@
which is independent on the choice of coarse-graining (provided the coarse-graining spaces are orthogonal).
In the case where the bath is dissipative rather than solely causing dephasing, the work bound is given by [25] (see Apendix A)
@@
$\geq\Delta E_{S}+\frac{1}{\beta}(H\left(p_{n}\right)-S_{A}(t_{0}))-\frac{1}{\beta}\Delta S_{SA}.$
The formal difference between the work bounds in Eq. 11 and Eq. 12 is $-\frac{1}{\beta}\Delta S_{SA}$. The term $\Delta S_{SA}$ is non-negative ($\Delta S_{SA}\geq 0$) because the unitary operation on the system and ancilla does not alter the system’s entropy, and the dephasing process is equivalent to an unread measurement. It is well established that entropy can only increase or remain constant during unread measurements [33]. Thus, the dissipation work bound is always less than or equal to the dephasing work bound. This is a first result of our analysis, which goes against the intuition that heat dissipation in the measuring apparatus is detrimental to the energy cost. The reason for that is the possibility in the case of a thermal bath to perform a reversible driving protocol in presence of the bath, achieving simultaneously the steps (1& 2) of the protocol. When resetting the apparatus, reversing the driving protocol allows for getting the invested work back entirely in the limit of quasi-static operation [25]. In contrast, pure-dephasing remains strictly irreversible and no meaningful reversible quasi-static limit can be obtained in general.
@@ IV Specific model: a qubit in a cavity
To push further our analysis, it is useful to examine a specific example of the protocol from Fig. 1. We consider that the system is a qubit (e.g. a two-level atom) and the ancilla is a harmonic oscillator (e.g. a cavity). After a suitable qubit-oscillator interaction, measuring excitations in the oscillator reveals the qubit’s energy state. This atom-cavity setup allows for the exploration of the impact of measurement strengths, from weak to strong. A key feature of this system, common in the case of weak measurements [28], is the large number of possible outcomes, determined by the ancilla’s infinite-dimensional space, compared to the qubit observables’ two eigenstates.
@@
$\hat{H}_{S,A}=\underbrace{\frac{\omega}{2}\hat{\sigma}_{z}\otimes\hat{I}}_{\hat{H}_{S}}+\underbrace{\hat{I}\otimes\omega_{a}\hat{a}^{\dagger}\hat{a}}_{\hat{H}_{A}}+\underbrace{\mu\left(t\right)\hat{\Pi}_{g}\otimes\left(\hat{a}-\hat{a}^{\dagger}\right)}_{\hat{V}_{SA}}.$ (13)
We have introduced the energy eigenbasis $\{{\left|{e}\right\rangle},{\left|{g}\right\rangle}\}$ of the qubit, the Pauli matrix $\hat{\sigma}_{z}={\left|{e}\right\rangle}{\left\langle{e}\right|}-{\left|{g}\right\rangle}{\left\langle{g}\right|}$ which corresponds to the measured system observable, and the ancilla’s bosonic anihilation operator $\hat{a}$. Over all this paper, we take $\hbar=1$. The coupling operator $\hat{V}_{SA}$ induces a coherent phase-space displacement of the cavity state if the qubit is in state ${\left|{g}\right\rangle}$. We assume that the function $\mu\left(t\right)$ has support only on the time interval $(t_{0},t_{1})$, outside which the coupling is completely switched off.
Before the pure dephasing occurs, we allow for an additional direct driving of the ancilla so as to enhance the measurement performance. We restrict our analysis to operations resulting in an unconditional phase-space displacement. The resulting unitary evolution then takes the form of a displacement operator
$\hat{D}\left(\alpha\right)=\exp\left(\alpha\hat{a}^{\dagger}+\alpha^{*}\hat{a}\right).$ (14)
Next, the ancilla is put in contact with a dephasing environment until it reaches a completely incoherent state in the number basis, at time $t_{\text{f}}$. The qubit-ancilla state then as the form (see Appendix B):
@@
where $\phi$ denotes the total phase accumulated during the interaction and driving processes and $\alpha_{1,2}$ the net displacement acquired by the cavity, conditioned to the qubit being in the ground and excited states, respectively. We have denoted $p_{j0}$, $j=e,g$ and $\rho_{ij0}$, $i,j=e,g$ the initial populations and coherences of the qubit.
Finally, the ancilla excitation number is read out, and the information is stored in a classical memory. The conditional state corresponding to the outcome $n$ is
$p_{g|n}\left(t{{}_{f}}\right)$ $=$ $\frac{p_{g0}}{p_{n}}e^{-|\alpha_{1}|^{2}}\frac{|\alpha_{1}|^{2n}}{n!}$ (16)
$\rho_{ge|n}\left(t{{}_{f}}\right)$ $=$ $\frac{\rho_{{ge}0}}{p_{n}}e^{\frac{-\left(|\alpha_{1}|^{2}+|\alpha_{2}|^{2}\right)}{2}}e^{-i{\phi}}\frac{\alpha_{1}^{n}}{\sqrt{n!}}\frac{\alpha_{2}^{*n}}{\sqrt{n!}}$
$\rho_{eg|n}\left(t{{}_{f}}\right)$ $=$ $\frac{\rho_{{eg}0}}{p_{n}}e^{\frac{-\left(|\alpha_{1}|^{2}+|\alpha_{2}|^{2}\right)}{2}}e^{i{\phi}}\frac{\alpha_{1}^{*n}}{\sqrt{n!}}\frac{\alpha_{2}^{n}}{\sqrt{n!}}$
$p_{e|n}\left(t{{}_{f}}\right)$ $=$ $\frac{p_{e0}}{p_{n}}e^{-|\alpha_{2}|^{2}}\frac{|\alpha_{2}|^{2n}}{n!},$
@@
The state update associated with finding $n$ excitations in the ancilla can be formulated in term of a measurement operator $\hat{M}_{n}$, according to $\hat{\rho}_{S|n}=\hat{M}_{n}\hat{\rho}_{S}\hat{M}_{n}^{\dagger}/p_{n}$, where:
@@
[…]
where $s_{r}$ represents the interval of values for $n$ corresponding to result $r$.
We are especially interested in achieving the limit of a weak measurement, which is obtained when the two displacements are assumed to be nearly identical. It is therefore convenient to re-express the measurement parameters as:
@@
V Figure of merits for the measurement performance
In a typical measurement, information is transferred from the system to the ancilla where it can be read. However, this transfer can be incomplete or noisy: a weak (incomplete) measurement leaves information unextracted in the system, while an inefficient measurement is subject to classical noise, or stores part of the extracted information in inaccessible degrees of freedom, resulting in information loss during the ancilla’s readout (see Fig. 2). While no standard measures exist to grasp those concepts in the full generality of the zoology of quantum measurements [5], we build below figures of merit to characterize the performance of the measurements captured by our model.
@@ V Figure of merits for the measurement performance
$\xi$ $=$ $\frac{S\left[\hat{\rho}_{S}(t_{f})\right]}{\log{d}},$ (20)
where it is understood that $\hat{\rho}_{S}(t)=\sum_{r}p_{r}{\cal E}_{r}[{\left|{\text{ref}}\right\rangle}{\left\langle{\text{ref}}\right|}]$, and $S\left(\hat{\rho}\right)$ denotes the von Neumann entropy of the density operator $\hat{\rho}$. We have normalized the strength measure $\xi$ so that $\xi=0$ means an absence of extracted information, while $\xi=1$ indicates maximal extraction of the observable’s statistics, leaving the system in a completely mixed state of the eigenstates ${\left|{j}\right\rangle}$. When $\xi\ll 1$, the measurement is weak; at $\xi=1$, it is strong, with intermediate values providing a scale to rank general measurements.
@@
This quantity generalizes the quantum efficiency of detectors appearing in weak continuous measurement scenarios [28, 35]. It is equal to 00 (resp. $1$) when none of the information (resp. all the information) extracted about the system is obtained by reading the ancilla. In the absence of coarse-graining (that is, $p(r|n)=\delta_{r,n}$), $\chi$ is equal to the quantum mutual information $I_{q}$ between the ancilla and the qubit at time $t_{f}$. In the presence of coarse-graining, $\chi\leq I_{q}$. However, in any case, $\chi$ is an upper bound on the correlations between the measurement outcome distribution $p(r)$ and the statistics of any observable of $S$ [34] (see also Appendix. E).
While the efficiency measures does quantify the ability of our measurement to actually make information about the system available in the ancilla, it is not always sufficient to eliminate bad quality measurements. Indeed, as illustrated in the next section, there are cases where the efficiency and the strength are maximal $(\xi=1,\eta=1)$, meaning that maximal correlations have built between the system and the accessible degrees of freedom of the ancilla, while the measurement result actually does not bring any information about the target measurement observable, or even about any system observable at all. While this occurs in somehow peculiar situations, where, e.g. the conditional states $\hat{\rho}_{S|r}$ are identical up to a phase in the measurement-eigenbasis coherences, it illustrates the necessity of another figure of merit. We use the classical mutual information $I\left(\{j\};\{r\}\right)=\sum_{j,r}p_{j,r}\log(p_{j,r}/p_{j}p_{r})$ between the classical output of the measurement, the result $r$, and the target observable statistics, initially by distribution $p_{j}={\left\langle{j}\right|}\hat{\rho}_{S}(t_{f}){\left|{j}\right\rangle}$. It is computed from the joint distribution $p_{j,r}={\left\langle{j}\right|}\hat{\rho}_{S|r}{\left|{j}\right\rangle}$ and $p_{r}=\sum_{n}p_{n}p(r|n)=\text{Tr}\{{\cal E}_{r}[\hat{\rho}_{S}(0)]\}$. The Holevo bound [36, 34] ensures that $I\left(\{j\};\{r\}\right)\leq\chi$ (see Appendix. E, such that a natural normalization for our third figure of merit $\eta_{X:r}$ is:
@@
This quantity vanishes when the measurement result $r$ and the observable value $x_{j}$ are two independent random variables; it reaches $1$ when all the system-ancilla correlations are all accounted for by the classical correlations in $p_{j,r}$. It is also worth mentioning that, unlike the efficiency measure $\eta$, the quantity $\eta_{X:r}$ is not independent on the measurement strength, and this dependence cannot be simply removed by renormalization. Together, these three figures of merit allow us to compare the performance of nonideal measurement of a system observable, as we do in the following for our qubit-cavity model. We finish by two remarks. First, owing to the chosen normalizations, the product of the three figures of merit verifies
$\xi\eta\eta_{X:r}=\frac{I\left(\{j\};\{r\}\right)}{\log{d}},$ (23)
@@
[…]

### v3

@@ III Thermodynamic analysis
We assume that the ancilla is at time $t_{0}$ in a free state at equilibrium with its environment at inverse temperature $\beta$.
@@
$\text{W}_{\text{dr}}=E_{SA}(t_{1})-E_{SA}(t_{0})=\Delta E_{S}+E_{A}(t_{1})-E_{A}(t_{0}),$ (5)
[…]
@@
We first analyze the case where the pure dephasing occurs in the eigenbasis of $\hat{H}_{A}$. During the dephasing process, there is no exchange of energy — neither work nor heat — with the systems. In the strong coupling regime, a nonzero work cost may be associated to the action of coupling and decoupling the system from the environment [mitchison]. However, the latter is not fundamental as it vanishes in the case where the coupling Hamiltonian to the environment commutes with the environment’s Hamiltonian (see Appendix F.). This is verified e.g. in the dispersive regime where the environment modes are far from resonance with the ancilla. We therefore exclude this cost from our lower bound. The second law then simply takes the form:
$\sigma=\Delta S_{SAM}\geq 0,$ (6)
where $\sigma$ denotes the entropy produced during the steps (2&3)(2\&3) and $\Delta S_{SAM}$ the change in von Neumann entropy of the system–ancilla–memory composite between the initial state and the state after the dephasing procedure. Additionally, the classical readout of the ancilla can, in principle, be implemented at no work cost [landauer1961irreversibility].
To provide a bound for the case where pure dephasing occurs in another basis, we consider that the Hamiltonian of the ancilla is switched to another Hamiltonian $\tilde{H}_{A}$ before the pure dephasing process starts. The latter is then assumed to damp coherences in the eigenbasis of $\tilde{H}_{A}$. Finally, the Hamilltonian of $A$ is switched back to $H_{A}$. Assuming instantaneous quench-like variations of the Hamiltonian, we identify an additional work cost [Esposito10]:
$W_{\text{sw}}$ $=$ $\text{Tr}\{(\tilde{H}_{A}-H_{A})(\hat{\rho}_{A}(t_{1})-\hat{\rho}_{A}(t_{f})\}$ (7)
$=$ $E_{A}(t_{1})-E_{A}(t_{f}).$
In the last line, we have used that the pure dephasing process conserves the expectation value of $\tilde{H}_{A}$. As before, the energy of $A$ is defined with respect to Hamiltonian $H_{A}$.
@@
$W^{A}_{\text{reset}}\geq{\color[rgb]{0,0,0}{E_{A}(t_{0})-E_{A}(t_{f})}}+\frac{1}{\beta}\Delta{\left\langle{S_{A}}\right\rangle},$ (8)
@@
represents the average change in entropy of the ancilla, realization-wise, expressed in terms of $\hat{\rho}_{A|r}=\text{Tr}_{S}\{\hat{\rho}_{SA|r}\}$. Based on the block-diagonal structure of $\hat{\rho}_{SA|r}$ in Eq. 4, the average entropy of the conditional ancilla states fulfills [nielsen2010quantum]:
$\sum_{r}p_{r}S[\hat{\rho}_{A|r}(t_{f})]=H\left(p_{n}\right)+\sum_{n}p_{n}H\left(p_{r|n}\right)-H\left(p_{r}\right),$ (10)
where $H\left(p_{i}\right)$ denotes the Shanon entropy of the probability distribution $p_{i}$.
@@
On the other hand, Landauer’s principle states that the work required to reset the classical memory fulfills [landauer1961irreversibility]:
$W_{\text{reset}}^{M}\geq\frac{1}{\beta}H\left(p_{r}\right).$ (11)
@@
which is independent on the choice of coarse-graining (provided the coarse-graining spaces are orthogonal), and of the pure-dephasing basis.
If we start with a pure initial ancilla state, it is straightforward to see that the work bound is bounded by $\log d_{A}$, where $d_{A}$ is the dimension of the Hilbert space of $A$. Therefore, a lower value is to be expected for small ancilla systems such as qubits than for higher-dimensional systems like a harmonic oscillator. However, we recall that in our description, the ancilla must be able to achieve macroscopically distinguishable pointer states, as does the amplified output of a measuring apparatus. It must therefore be a many-level system that is read out in a coarse-grained fashion (hence a typically small number of different outcomes). In the next section, we model the pointer as a harmonic oscillator so as to explore the consequences of this coarse-graining on the tradeoff between measurement cost and efficiency.
In the case where the bath is dissipative rather than solely causing dephasing, the work bound is given by [latune2024thermodynamically] (see Apendix A)
@@
$\geq\Delta E_{S}+\frac{1}{\beta}(H\left(p_{n}\right)-S_{A}(t_{0}))-\frac{1}{\beta}\Delta S_{SA},$
where $\Delta S_{SA}$ is the variation of the system and ancilla entropy due to the dephasing process, which is necessary nonnegative, as entropy can only increase or remain constant during unread measurements [nielsen2010quantum]. The formal difference between the work bounds in Eq. 12 and Eq. 13 is therefore $\frac{1}{\beta}\Delta S_{SA}$, such that the dissipation work bound is always less than or equal to the dephasing work bound. This is a first result of our analysis, which goes against the intuition that heat dissipation in the measuring apparatus is detrimental to the energy cost. The reason for that is the possibility in the case of a thermal bath to perform a reversible driving protocol in the presence of the bath, achieving simultaneously the steps (1& 2) of the protocol. When resetting the apparatus, reversing the driving protocol allows for getting the invested work back entirely in the limit of quasi-static operation [latune2024thermodynamically]. In contrast, pure-dephasing remains strictly irreversible and no meaningful reversible quasi-static limit can be obtained in general.
@@ IV Specific model: a qubit in a cavity
To push further our analysis, it is useful to examine a specific example of the protocol from Fig. 1. We consider that the system is a qubit (e.g. a two-level atom) and the ancilla is a harmonic oscillator (e.g. a cavity). After a suitable qubit-oscillator interaction, measuring excitations in the oscillator reveals the qubit’s energy state. This atom-cavity setup allows for the exploration of the impact of measurement strengths, from weak to strong. A key feature of this system, common in the case of weak measurements [jacobs2006straightforward], is the large number of possible outcomes, determined by the ancilla’s infinite-dimensional space, compared to the qubit observables’ two eigenstates.
@@
$\hat{H}_{S,A}=\underbrace{\frac{\omega}{2}\hat{\sigma}_{z}\otimes\hat{I}}_{\hat{H}_{S}}+\underbrace{\hat{I}\otimes\omega_{a}\hat{a}^{\dagger}\hat{a}}_{\hat{H}_{A}}+\underbrace{\mu\left(t\right)\hat{\Pi}_{g}\otimes\left(\hat{a}-\hat{a}^{\dagger}\right)}_{\hat{V}_{SA}},$ (14)
where $\omega_{a}$ is the harmonic oscillator frequency and $\omega$ the qubit frequency. We have introduced the energy eigenbasis $\{{\left|{e}\right\rangle},{\left|{g}\right\rangle}\}$ of the qubit, the Pauli matrix $\hat{\sigma}_{z}={\left|{e}\right\rangle}{\left\langle{e}\right|}-{\left|{g}\right\rangle}{\left\langle{g}\right|}$ which corresponds to the measured system observable, and the ancilla’s bosonic anihilation (creation) operator $\hat{a}$ $\left(\hat{a}^{\dagger}\right)$. Over all this article, we take $\hbar=1$. The coupling operator $\hat{V}_{SA}$ induces a coherent phase-space displacement of the cavity state if the qubit is in state $\left|{g}\right\rangle$}. We assume that the function $\mu\left(t\right)$ has support only in the time interval $(t_{0},t_{1})$, outside which the coupling is completely switched off.
To explore a broader class of measurements, we allow for an additional direct driving of the ancilla before the pure dephasing occurs. We restrict our analysis to operations resulting in an unconditional phase-space displacement. The resulting unitary evolution then takes the form of a displacement operator
[…]
Next, the ancilla is put in contact with dephasing environments until it reaches a completely incoherent state in the number basis, at time $t_{\text{f}}$. The qubit-ancilla state then has the form (see Appendix B):
@@
where $\phi$ denotes the total phase accumulated during the interaction and driving processes and $\alpha_{1,2}$ the net displacement acquired by the cavity, conditioned on the qubit being in the ground and excited states, respectively. We have denoted $p_{j0}$, $j=e,g$ and $\rho_{ij0}$, $i,j=e,g$ the initial populations and coherences of the qubit. The cavity displacements $\alpha_{1}$ and $\alpha_{2}$ associated with states $\left|{g}\right\rangle$} and $\left|{e}\right\rangle$}, respectively, are related to the conditional displacement $\alpha_{\text{cond}}$ generated by the Hamiltonian Eq. (14) and the second displacement $\alpha$ via $\alpha_{1}=\alpha_{\text{cond}}+\alpha$ and $\alpha_{2}=\alpha$. Finally, the ancilla excitation number is read out, and the information is stored in a classical memory.
The conditional state corresponding to the outcome $n$ is
[…]
@@
The state update associated with finding $n$ excitations in the ancilla can be formulated in terms of a measurement operator $\hat{M}_{n}$, according to $\hat{\rho}_{S|n}=\hat{M}_{n}\hat{\rho}_{S}\hat{M}_{n}^{\dagger}/p_{n}$, where:
@@
$\hat{\rho}_{S|r}=\mathcal{E}_{r}\left[\hat{\rho}_{S}(0)\right]=\sum_{n\hskip 0.85358pt\in\hskip 0.85358pts_{r}}\hat{M}_{n}\hat{\rho}_{S}(0)\hat{M}^{\dagger}_{n},$ (19)
where $s_{r}$ represents the interval of values for $n$ corresponding to the result $r$.
We are particularly interested in the weak measurement limit, attained when the final displacements applied to the ground and excited states are nearly identical. It is therefore convenient to re-express the measurement parameters as:
@@
V Figures of merit for the measurement performance
In a typical measurement, information is transferred from the system to the ancilla where it can be read. However, this transfer can be incomplete or noisy: a weak (incomplete) measurement leaves information unextracted in the system, while an inefficient measurement is subject to classical noise, or stores part of the extracted information in inaccessible degrees of freedom, resulting in information loss during the ancilla’s readout (see Fig. 3). While no standard measures exist to grasp those concepts in the full generality of the zoology of quantum measurements [wiseman2009quantum], we build below figures of merit to characterize the performance of the measurements captured by our model.
@@ V Figures of merit for the measurement performance
We first propose a measure for the strength of the measurement associated with the total information extracted from the system. Starting from our pure reference state $\left|{\text{ref}}\right\rangle$}, the latter can be quantified by the von Neumann entropy of the unconditioned output state of the system $\hat{\rho}_{S}(t_{f})=\sum_{r}p_{r}\hat{\rho}_{S|r}=\text{Tr}_{A}\hat{\rho}_{SA}(t_{f})$. “Unconditioned” here means averaged over all measurement outcomes, which corresponds to the case where the measurement outcomes are either inaccessible or simply not registered in the classical memory. Indeed, the purity decrease of the average system state is entirely due to correlations built between the system and other degrees of freedom (including the ancilla and its environment) such that it quantifies (or at least upper bounds) the information about the system that can be acquired by measuring those other degrees of freedom. We stress that the von Neumann entropy variation during the measurement depends both on the measurement performance and on the system’s initial state (e.g. if it is already in a maximally mixed state, the entropy does not increase further); hence, the necessity of a reference state. In summary, we quantify the strength of the measurement via:
@@
The measurement performance is also limited by the information “lost” in the environment (that is, which cannot be retrieved in the accessible degrees of freedom of the ancilla, but could be read in principle in other degrees of freedom of the ancilla or in the environment). Still considering that the system starts in the reference state $\left|{\text{ref}}\right\rangle$}, this loss can be quantified from the purity (or von Neumann entropy) of the conditional system state $\rho_{S|r}=\text{Tr}\{\rho_{SA|r}\}$ associated with obtaining the result $r$. Indeed, a pure conditional state means that reading the ancilla allowed the observer to obtain all the information transferred during the system-ancilla interaction, while a mixed conditional state means that more information could be obtained by reading the inaccessible degrees of freedom. To obtain a quantity independent of both the measurement outcome and the measurement strength, and maximal when all available information is accessed, we use the Holevo information [Wilde2011Jun] $\chi=S[\hat{\rho}_{S}]-\sum_{r}p_{r}S(\hat{\rho}_{S|r})\geq 0$. which takes its maximum value $S[\hat{\rho}_{S}]$ when all the extracted information about $S$ can be read in the ancilla. Normalizing to this maximum value leads to the following figure of merit for the measurement efficiency:
@@
This quantity generalizes the notion of quantum efficiency for detectors in weak continuous measurement scenarios [jacobs2006straightforward, Wiseman1993Jan, ficheux2018dynamics], and can be identified from the stochastic master equations governing the dynamics under continuous measurement.
The efficiency $\eta$ is equal to 0 (resp. $1$) when none of the information (resp. all the information) extracted about the system is obtained by reading the ancilla. In the absence of coarse-graining (that is, $p(r|n)=\delta_{r,n}$), $\chi$ is equal to the quantum mutual information $I_{q}$ between the ancilla and the qubit at time $t_{f}$. In the presence of coarse-graining, $\chi\leq I_{q}$. However, in any case, $\chi$ is an upper bound on the correlations between the measurement outcome distribution $p(r)$ and the statistics of any observable of $S$ [Wilde2011Jun] (see also Appendix. E).
While the efficiency measure does quantify the ability of our measurement to actually make information about the system available in the ancilla, it is not always sufficient to eliminate bad quality measurements.
Indeed, as illustrated in the next section, there are cases where the efficiency and the strength are maximal $(\xi=1,\eta=1)$, meaning that maximal correlations have built between the system and the accessible degrees of freedom of the ancilla, and yet the measurement result actually does not bring any information about the target measurement observable, or even about any system observable at all. This effects occurs whenever the ancilla is not read in the optimal basis, as noted e.g. in superconducting qubit weak measurement experiments [murch2013observing].
While this issue is avoided in well-calibrated measurement setups, it illustrates the necessity an additional figure of merit in order to fully characterize the quality of arbitrary measurements. We use the classical mutual information $I\left(\{j\};\{r\}\right)=\sum_{j,r}p_{j,r}\log(p_{j,r}/p_{j}p_{r})$ between the classical output of the measurement, the result $r$, and the target observable statistics, initially by distribution $p_{j}={\left\langle{j}\right|}\hat{\rho}_{S}(t_{f}){\left|{j}\right\rangle}$. It is computed from the joint distribution $p_{j,r}={\left\langle{j}\right|}\hat{\rho}_{S|r}{\left|{j}\right\rangle}$ and $p_{r}=\sum_{n}p_{n}p(r|n)=\text{Tr}\{{\cal E}_{r}[\hat{\rho}_{S}(0)]\}$. The Holevo bound [Preskill16, Wilde2011Jun] ensures that $I\left(\{j\};\{r\}\right)\leq\chi$ (see Appendix. E, such that a natural normalization for our third figure of merit $\eta_{X:r}$ is:
@@
This quantity vanishes when the measurement result $r$ and the observable value $x_{j}$ are two independent random variables; it reaches $1$ when all the system-ancilla correlations are accounted for by the classical correlations in $p_{j,r}$.
It is also worth mentioning that, unlike the efficiency measure $\eta$, the quantity $\eta_{X:r}$ is not independent of the measurement strength, and this dependence cannot be simply removed by renormalization. $\eta_{X:r}$ is therefore not equivalent to $\eta$, and rather brings additional information about the measurement efficiency. Together, these three figures of merit allow us to analyze the performance of nonideal measurements of a system observable, as we do in the following for our qubit-cavity model.
We finish by two remarks. First, owing to the chosen normalizations, the product of the three figures of merit verifies
[…]
@@
Figure 3: Hierarchy of information flows involved in the figures of merit. The inequality from Eq. 25 suggests the following interpretation: From the total amount of information $\log d$ missing to characterize the value of $\hat{X}$ in the state $\left|{\text{ref}}\right\rangle$}, a fraction $S[\hat{\rho}_{S}(t_{f})]$ is transferred to the ancilla and environment (mustard arrow). A smaller amount $I_{q}$ is available in the ancilla (pink arrow), and an even smaller amount $\chi$ in the practically accessible degrees of freedom of the ancilla (grey arrow). Finally, only an amount I(j;r)I({j};{r}) concerns the target observable (blue arrow).